3. a. Briefly explain Ohm’s Law (2 marks)

b. A 750 Watt electric convector heater is plugged into a 240 Volt power supply. A trolley

passing over the flex from the heater to the plug has caused the insulation to distort

and has reduced the resistance between the live and neutral wires to 60 Ohms.

(i) What is the rate of heat production (in Watts) at the point where the cable is

damaged? (Please show your calculations.) (1 mark)

(ii) The plug fitted to the heater flex carries a l3 amp fuse, is this the correct rating

for the appliance? (2 marks)

(iii) Would the correctly rated fuse have 'blown' as a result of the fault? (Please

show your calculations.) (1 mark)

b. Several years ago a change in the recognised colour conventions led to phase (live)

cables in industrial three phase power installations being changed from red, yellow

and blue to black, grey and brown. What are the implications for premises that have

not been rewired since that time? (4 marks)

c. Electric lighting is often overlooked as a potential source of ignition. List factors that

might lead to an electric light fitting becoming a source of ignition. (5 marks)

d. List the problems associated with the use of electric adaptors in the workplace and

why their use should be minimised. (5 marks



1
Expert's answer
2021-02-02T09:30:34-0500

3. a. Briefly explain Ohm’s Law: current increases when voltage increases and current decreases when resistance increases.

b. (i) The rate of heat production is


"q=I^2R=\frac{V^2}{R}=960\text{ W}."

(ii) I3 amp fuse is for 1 A, the current is 750/240=3.1 A, this is a wrong fuse.

(iii) Yes, the current would be 240/60=4 A.

b. During the following electrical works in the building electricians need to use wires of accepted colors.

c. Something (like fabric/paper) left on a lamp or below the lamp.

d. At a workplace, workers have limited space and move their limbs a lot, which increases the risk of electrical shock due to unintended touch.